# A Survey of AI Agent Registry Solutions

### v1

@@ Abstract
As autonomous AI agents scale across cloud, enterprise, and decentralized environments, the need for standardized registry systems to support discovery, identity, and capability sharing has become essential. This paper surveys three prominent registry approaches each defined by a unique metadata model: MCP’s mcp.json, A2A’s Agent Card, and NANDA’s AgentFacts. MCP uses a centralized metaregistry with GitHub-authenticated publishing and structured metadata for server discovery. A2A enables decentralized interaction via JSON-based Agent Cards, discoverable through well-known URIs, curated catalogs, or direct configuration. NANDA Index introduces AgentFacts, a cryptographically verifiable and privacy-preserving metadata model designed for dynamic discovery, credentialed capabilities, and cross-domain interoperability.These approaches are compared across four dimensions: security, scalability, authentication, and maintainability.The paper concludes with suggestions and recommendations to guide future design and adoption of registry systems for the Internet of AI Agents.
@@ 1 Introduction
This paper presents a comparative analysis of four registry architectures, each offering a distinct approach to addressing these challenges:
  • MCP Registry (Anthropic, 2025) [1] which provides a centralized metadata layer using structured mcp.json files for agent discovery and installation .
  • A2A Agent Cards [2] which describe agent capabilities and endpoints in a standardized JSON format, enabling discovery via well-known URLs, curated catalogs, or private distribution.
  • Microsoft Entra Agent ID (Microsoft, 2025) [3]: a SaaS-delivered, enterprise-grade directory built into Azure AD, offering built-in lifecycle, governance, and zero-trust controls.
  • NANDA Index: AgentFacts, (MIT AIDE, 2025)  [4] a decentralized, cryptographically verifiable format supporting dynamic resolution, credentialed assertions, and privacy-preserving queries.
@@ 3 MCP Registry
The MCP registry is a centralized “metaregistry” for discovering and installing MCP servers. Publishers push a versioned mcp.json via a CLI tool that performs a GitHub OAuth flow and, for reverse-DNS namespaces, a DNS TXT challenge. The Go-based REST API exposes read endpoints (no authentication) and write endpoints (GitHub OAuth + DNS verification), stores raw JSON in object storage, indexes metadata in MongoDB (with an in-memory fallback), and generates asynchronous jobs or webhooks. Downstream MCP client apps poll the registry (or private mirrors), cache the data locally, and serve end-users without direct live calls to the central service.
@@ 6 NANDA Index
    AgentFacts Layer: Distributes rich, schema-validated metadata using self-describing JSON-LD documents signed as W3C Verifiable Credentials. These documents describe capabilities, endpoints, telemetry, and authentication logic, and can be hosted at agent-controlled domains or decentralized storage (e.g., IPFS), allowing privacy-preserving updates independent of the index.
@@ 6.3 Scalability
The system supports federated deployment, where Lean Index instances can be domain-specific (e.g., healthcare, finance) or geo-localized. Cross-instance interoperability is achieved through DID-based resolution and verifiable claims, allowing seamless operation across federated trust boundaries.
@@ 6.4 Maintainability
Table 1: Comparison of AI Agent Registry solution by A2A, MCP and NANDA Index
Dimension MCP Google A2A NANDA Index
Purpose Enable publication and discovery of agent servers for LLM tool use Advertise a server-side agent’s endpoint and skills for JSON-RPC Convey live endpoints plus cryptographically-signed capabilities, SBOM hash, privacy path, revocation info
Discovery Path REST: /v0/servers, /v0/servers/{id} One-step fetch at /.well-known/agent.json or central catalogue Two-step: lean registry $\rightarrow$ FactsURL / PrivateFactsURL
Trust Primitive Bearer token + GitHub credentials Plain HTTPS and optional token, self-declared attributes W3C Verifiable Credential v2 signatures; VC-Status revocation (<1 s), support for third-party audited attributes
Privacy Option Container metadata and manifest publishing; optional offline operation None (lookup hits agent host) Optional PrivateFactsURL via IPFS/Tor; hides requester
Endpoint Freshness Timestamps on publish: created_at, updated_at; dynamic via REST Assumes minutes-to-hours stability (no TTL field) TTL-scoped lists: static (1–6 h) or rotating (5–15 m)
Schema Weight Approx. 1–3 KB JSON $\approx$ 0.3–1 KB JSON 1–3 KB JSON-LD + VC
Best-fit Use Case Centralized registry of agent/server tools (Docker/NPM); designed for scalable deployment and LLM plug-in integration Stable SaaS agents inside a single marketplace Highly mobile, privacy-sensitive, or safety-critical agents at trillion scale
Table 2: Detailed Feature Comparison: Agent Card vs. Agent Facts
Feature Agent Card (Google A2A) Agent Facts (NANDA Index)
Identifier url (host-tied) Id
Label & description name, description, version agent name, label, description, version
Provider Info provider.organization, url provider.{name,url,id}
Static endpoint single url "endpoints.static"[ ]
Rotating / geo endpoints "endpoints.rotating"[ ] (URL + TTL)
Adapter router "endpoints.adaptive_router".url
Skills / Capabilities skills[ ], capabilities.* skills[ ] + "capabilities"
Auth Schemes securitySchemes, security capabilities.authentication
Protocol Flags capabilities.streaming, pushNotifications capabilities.streaming
SBOM/ Integrity sbomHash (in VC)
Telemetry telemetry.{endpoint,sampling}
Performance/audit evaluations block
Privacy Path private_facts_url
Cryptographic wrapper plain JSON, HTTPS JSON-LD + W3C VC
optional signature
TTL/Freshness none (HTTP cache) per-endpoint TTL fields
Size 0.3 – 1 KB 1–3 KB
Fetch hops 1 (well-known) 2 (registry $\rightarrow$ Facts)
Revocation speed HTTP re-fetch (min) VC-Status (<1 s)
Implementation effort low higher (JSON-LD, VC Toolchain)
@@ 7 Comparative Evaluation
  • Table 1 compares registry solutions from MCP, A2A (Google), and NANDA Index, analyzing their purpose, discovery paths, trust primitives, privacy mechanisms, endpoint freshness strategies, schema complexity, and best-fit use cases.
@@
8 Phase-wise Evolution of Agent Registry Architectures
@@
8.1 Phase I — Static, Isolated Discovery
@@
8.2 Phase II — Dynamic RESTfull APIs
@@
8.3 Phase III — Verifiable Metadata and Federated Trust and AI agent quilt
@@
9 Conclusion
The proliferation of autonomous AI agents across enterprise, research, and consumer domains has created an urgent infrastructure challenge: how to discover, identify, and trust agents at Internet scale. This survey examined four representative registry architectures: MCP Metaregistry, A2A, Microsoft Entra Agent ID and NANDA Index, each addressing core requirements of security, authentication, scalability, and maintenance through distinct architectural approaches.
@@ 9 Conclusion
    Decentralization Enables Long-term Sustainability. While centralized approaches offer operational simplicity, they create single points of failure and vendor lock-in risks that become increasingly problematic as agent ecosystems mature. NANDA Index’s federated design demonstrate how decentralized architectures can achieve both scalability and community governance. The dual-path resolution pattern in NANDA particularly addresses privacy concerns that will become critical as agent interactions proliferate.
@@
    Security Must Be Built-in, Not Bolted-on. All examined registries recognize that cryptographic integrity is foundational whether through W3C Verifiable Credentials (NANDA), DNS-TXT verification (MCP), or Azure AD’s enterprise security controls (Entra). However, the separation of static identity resolution from dynamic metadata distribution, pioneered by NANDA, offers the most robust foundation for preventing registry poisoning and enabling privacy-preserving discovery patterns.

### v2

# Evolution of AI Agent Registry Solutions: Centralized, Enterprise, and Distributed Approaches

@@ Abstract
Autonomous AI agents now operate across cloud, enterprise, and decentralized domains, creating demand for registry infrastructures that enable trustworthy discovery, capability negotiation, and identity assurance. We analyze five prominent approaches: (1) MCP Registry (centralized publication of mcp.json descriptors), (2) A2A Agent Cards (decentralized self-describing JSON capability manifests), (3) AGNTCY Agent Directory Service (IPFS Kademlia DHT content routing extended for semantic taxonomy–based content discovery, OCI artifact storage, and Sigstore-backed integrity), (4) Microsoft Entra Agent ID (enterprise SaaS directory with policy and zero-trust integration), and (5) NANDA Index AgentFacts (cryptographically verifiable, privacy-preserving fact model with credentialed assertions). Using four evaluation dimensions—security, authentication, scalability, and maintainability—we surface architectural trade-offs between centralized control, enterprise governance, and distributed resilience. We conclude with design recommendations for an emerging Internet of AI Agents requiring verifiable identity, adaptive discovery flows, and interoperable capability semantics.
@@ 1 Introduction
This paper presents a comparative analysis of five registry architectures:
  • MCP Registry [1]: centralized metadata publication for MCP servers via versioned mcp.json.
  • A2A Agent Cards [2]: decentralized JSON capability and endpoint description resolved at /.well-known/agent.json.
  • AGNTCY Agent Directory Service (ADS): IPFS Kademlia DHT content routing (extended for semantic taxonomy discovery), OCI artifact storage, Sigstore integrity proofs, and enterprise federation hooks.
  • Microsoft Entra Agent ID [3]: managed enterprise directory integrating lifecycle, governance, conditional access, and policy enforcement.
  • NANDA Index: AgentFacts [4]: decentralized verifiable credential (VC)-backed fact model supporting privacy-preserving dual-path discovery.
@@ 2 Background
  • AGNTCY Agent Directory Service (ADS): A content-addressed, OCI-aligned directory that resolves semantic capabilities to immutable digests and uses decentralized rendezvous (DHT) for provider discovery [6].
@@ 3 MCP Registry
The MCP registry is a centralized “metaregistry” for discovering and installing MCP servers. Publishers push a versioned mcp.json via a CLI tool that performs a GitHub OAuth flow and, for reverse‐DNS namespaces, a DNS TXT challenge. The Go-based REST API exposes read endpoints (no authentication) and write endpoints (GitHub OAuth + DNS verification), stores raw JSON in object storage, indexes metadata in MongoDB (with an in-memory fallback), and generates asynchronous jobs or webhooks. Downstream MCP client apps poll the registry (or private mirrors), cache the data locally, and serve end-users without direct live calls to the central service.
@@ 4.4 Maintainability
5 AGNTCY Agent Directory Service
The AGNTCY Agent Directory Service (ADS) is organized around four composable layers that separate data modeling, storage, distribution, and content selection so the system can scale and evolve independently at each concern:
  1. 1.
    Agent AI metadata modeling (Open Agent Schema Framework, OASF). Open Agent Schema Framework (OASF) cleanly separates what an agent can do from the conditions under which it may do it. Semantic attributes describe the capability surface—skills and operational domains. Constraint descriptors capture required dependencies (models, tools, data sources), resource and latency cost profiles, token budgets, and any compliance gates. This separation lets a resolver align intent with capability while enforcing execution bounds.
  2. 2.
    Immutable object storage (OCI). Directory records are packaged as OCI-compliant artifacts and pushed to registries; each artifact is addressed by a cryptographic digest that enforces immutability and enables de-duplication. The OCI naming model yields globally unique references and efficient layer reuse for transport and caching. Human-readable aliases (tags or descriptive name attributes) can be attached without weakening the canonical digest binding.
  3. 3.
    Content distribution (OCI Distribution Spec). Global replication, caching, and mirroring leverage the standard OCI Distribution Specification [7], allowing any compliant registry to host identical artifacts (same digest) without coordination—enabling multi-source failover and locality-aware retrieval.
  4. 4.
    Content and registry selection (IPFS Kademlia DHT). Selection reconciles heterogeneous registries via a two–step mapping: (1) semantic taxonomies (skills, domains) resolve to canonical OCI content identifiers (digests) through capability→digest indices; (2) each digest is mapped to one or more registry identifiers and locators (endpoints, tags, replica metadata). The extended Kademlia DHT [8] stores and routes both mapping layers, letting the content scheduler first derive the immutable content address, then choose optimal registries/replicas using locality and freshness cues. Once a digest is resolved and its minimal record (plus referenced artifact metadata) retrieved, rich filtering across the full OASF attribute set occurs locally; the DHT acts only as a rendezvous layer optimized for sparse semantic capability mappings rather than exhaustive attribute indexing.
On Naming and Zooko’s Triangle.
By anchoring human-readable attributes to cryptographic OCI digests (via signed OASF records) and resolving them through a decentralized DHT, ADS satisfies the three traditionally conflicting properties of Zooko’s triangle: (1) decentralized naming, (2) secure naming (Sigstore provenance plus signed attestations [9]), and (3) human-readable naming (scoped semantic attributes and constraints). A common criticism is that the “secure” leg, when implemented solely with Sigstore, still relies on a partially centralized trust root. To mitigate this, ADS can augment Sigstore with a trustless, smart-contract–anchored attestation path (e.g., ERC-8004 [10]) that commits record signature digests on-chain.
Architecture
Building on the four layers above, ADS introduces a semantic enrichment path. Hierarchical skill taxonomies accelerate capability-scoped queries before DHT resolution. Replica metadata (latency, region, freshness) feeds a composite scoring function that produces deterministic, low-cost selection decisions. Location-independent OCI digests (sha256:<digest>) enable transparent multi-registry reconciliation: any registry serving the same digest becomes an interchangeable source. Digest equality, reinforced by Sigstore provenance attestations, enforces immutability and prevents substitution. Dynamic selection then avoids centralized choke points by coupling semantic relevance with Kademlia routing proximity.
Publisher Agent Directory API Local Search (Client) OCI Storage Global Search (DHT / Index) Content & Remote Selection Push artifact (digest)Submit minimal OASF recordSkill $\to$ digest mappingDigest $\to$ replica endpointsACK / statusPublish complete (a) Publishing Flow: artifact pushed first; minimal record + sparse semantic mapping + replica endpoints inserted. Capability / embedding queryCandidate digestsFetch minimal recordsResolve digest $\to$ endpointsRanked endpointsRecords + endpointsPull artifact (digest) (b) Discovery Flow: DHT/global search supplies only sparse skill $\to$ digest; rich filtering happens locally using retrieved minimal records.
Figure 1: Time-flow (sequence) diagrams for AGNTCY Agent Directory. Vertical lifelines represent: Publisher, Agent Directory API, Local Search, OCI Storage, Global Search (semantic DHT / embedding index), and Content & Remote Server Selection. (a) Publishing: artifact push precedes minimal record publication; sparse semantic (skill $\to$ digest) plus digest $\to$ endpoint mappings are inserted. (b) Discovery: local capability intent resolves via global search to digests, minimal records are fetched, local filtering/ranking occurs, endpoints resolved, and the selected artifact is pulled.
5.1 Security
ADS security separates (1) content / artifact security from (2) server / runtime security, allowing strong integrity and provenance guarantees even in minimally trusted network environments.
5.2 Authentication
Authentication in AGNTCY ADS is implemented in two distinct parts: data authenticity and server authentication.
Data Authenticity. This ensures that the data producer is authenticated and the integrity of the data is verifiable. All directory records are packaged as OCI records whose cryptographic digests are self-authenticating: the name (sha256:<digest>) is a direct function of the bytes, giving immutable, collision-resistant identifiers and secure naming by construction. Each published digest is accompanied by Sigstore provenance and signature materials (provenance attestation, record, producer identity, source revision), enabling pre-execution policy such as “only activate if provenance + record verified” without trusting a central admission service. Capability and constraint descriptors in the minimal OASF record are signed and bound (via the referenced digest) to the record they describe, preventing capability re-binding or privilege escalation through manifest substitution.
Server Authentication. This ensures trust among server peers and restricts unauthorized access. Core discovery can be left open (content trust is intrinsic), but operators often restrict which peers a directory node will accept or exchange routing/index data with. “Secure mode” introduces: (1) authenticated peer admission (static allowlist or dynamic attestation) before Kademlia session establishment; (2) mutual TLS with short-lived SPIFFE/SPIRE SVIDs or equivalent rotating credentials. These measures shrink the unsolicited attack surface (e.g., DDoS amplification/state exhaustion) and satisfy enterprise policies that forbid arbitrary federation.
5.3 Scalability
Every directory node contributes routing, semantic index shards, and OCI artifact cache capacity. Adaptive replication promotes hot capability shards and frequently requested artifact digests near demand clusters. IPFS Kademlia iterative lookups (extended with locality + semantic relevance scoring) are augmented by locality-aware peer selection (latency + load) while opportunistic caching shortens future paths. Joint optimization of OCI naming and distribution means that the same digest can be resolved to the lowest-latency verified replica; if replicas diverge the digest changes, forcing explicit revalidation. Content prefetch and layer deduplication (OCI) reduces bandwidth while preserving deterministic digest integrity.
5.4 Maintainability
Declarative deployment (Kubernetes with Helm) together with GitOps workflows minimizes control-plane drift. Cross-registry reconciliation through digest equality lets autonomous registries evolve naming and tagging conventions independently while remaining interoperable at the content-address level. Interoperability with external registries is a distinctive feature: data from sources such as MCP or A2A can be reconciled provided it is modeled into OASF via the corresponding module. Modules for MCP and A2A are currently available.
@@ 7 NANDA Index
    AgentFacts Layer: Distributes rich, schema-validated metadata using self-describing JSON-LD documents signed as W3C Verifiable Credentials. These documents describe capabilities, endpoints, and authentication logic, and can be hosted at agent-controlled domains or decentralized storage (e.g., IPFS), allowing privacy-preserving updates independent of the index.
@@ 7.3 Scalability
The system supports federated deployment, where Lean Index instances can be domain-specific (e.g., healthcare, finance) or geo-localized. Cross-instance interoperability is achieved through DID-based resolution and verifiable claims, allowing seamless operation across federated trust boundaries. For neutrality, the Index will also be hosted at 15 universities to enable resilience and distributed governance.
@@ 7.4 Maintainability
Table 1: Comparison of Five Agent Registry Approaches
Dimension MCP A2A AGNTCY Entra Agent ID NANDA Index
Purpose Centralized publish + discover for MCP servers Self-hosted capability + endpoint descriptor Distributed + multi-registry interoperability Managed enterprise agent directory Verifiable, privacy-preserving capability facts
Discovery Path REST list + GET by id Well-known JSON (1 hop) Semantic discovery based on taxonomies with local and global metasearch Portal + Graph/Policy APIs Lean index $\rightarrow$ Facts / PrivateFacts (2 hop)
Trust Primitive GitHub OAuth + DNS TXT HTTPS + optional token Sigstore for data provenance attestations, trustless EVM optional. Azure AD token + policy engine VC v2 signatures + VC-Status
Privacy Option None (public reads) None Notion of private directories and public vs public records. Directory scope policies PrivateFactsURL (obfuscated path)
Endpoint Freshness Poll + updated timestamps Assumed stable (no TTL) Adaptive DHT cache; digest rev Platform-managed sync TTL + rotating / adaptive endpoints
Schema Weight 1–3 KB JSON 0.3–1 KB JSON  4MB + signature Directory object metadata 1–3 KB JSON-LD + VC
Best Fit Tool/plugin ecosystems SaaS-style API agents Federated / hybrid fleets, multi-registry Regulated enterprise governance High-churn, privacy / verifiability critical
Table 2: Agent Advertisement Models: Card vs. Facts
Feature A2A Agent Card NANDA AgentFacts
Identifier Host URL Stable ID + signed AgentAddr ref
Endpoints Single fixed URL Static + rotating + adaptive (TTL) set
Capability Model skills + capabilities.* skills + capabilities (VC claims)
Authentication securitySchemes list capabilities.authentication + VC issuer attestations
Integrity Wrapper Plain JSON over HTTPS JSON-LD + VC signature + status list
Privacy Option None PrivateFactsURL (dual-path)
Freshness HTTP cache only TTL per endpoint + revocation events
Revocation Re-fetch interval VC-Status (<1s)
Size (typical) 0.3–1 KB 1–3 KB
Fetch Hops 1 (well-known) 2 (Index → Facts)
Implementation Effort Low Medium (VC tooling)
@@ 8 Comparative Evaluation
  • Table 1 compares registry solutions from MCP, A2A (Google), AGNTCY, Entra Agent ID and NANDA Index, analyzing their purpose, discovery paths, trust primitives, privacy mechanisms, endpoint freshness strategies, schema complexity, and best-fit use cases.
@@
9 Evolution of Agent Registry Architectures
@@
9.1 Static, Isolated Discovery
@@
9.2 Dynamic RESTfull APIs
@@
9.3 Verifiable Metadata and Federated Trust and AI agent quilt
@@
10 Discussion
The proliferation of autonomous AI agents across enterprise, research, and consumer domains has created an urgent infrastructure challenge: how to discover, identify, and trust agents at Internet scale. This survey examined five representative registry architectures: MCP Metaregistry, A2A, Microsoft Entra Agent ID, NANDA Index, and AGNTCY ADS, each addressing core requirements of security, authentication, scalability, and maintenance through distinct architectural approaches.
@@ 10 Discussion
    Decentralization Enables Long-term Sustainability. While centralized approaches offer operational simplicity, they create single points of failure and vendor lock-in risks that become increasingly problematic as agent ecosystems mature. NANDA Index’s federated design and AGNTCY ADS’s content-addressed, OCI-aligned model with decentralized rendezvous demonstrate how decentralized architectures can achieve both scalability and community governance. The dual-path resolution pattern in NANDA and ADS’s separation of semantic mapping from immutable content address both privacy and integrity concerns.
@@
    Security Must Be Built-in, Not Bolted-on. All examined registries recognize that cryptographic integrity is foundational whether through W3C Verifiable Credentials (NANDA), DNS-TXT verification (MCP), or Azure AD’s enterprise security controls (Entra). However, AGNTCY ADS stands out with its Sigstore-backed keyless signing and optional DID-based identity, ensuring both data authenticity and server authentication. This approach prevents registry poisoning and enables privacy-preserving discovery patterns.
@@
11 The Path Ahead: A Switchboard for the Agentic Web
Registry architectures must evolve toward federated models that separate stable identity resolution from dynamic capability metadata, enabling cryptographic trust and cross-organizational discovery without sacrificing privacy or operational autonomy: whether through protocol-specific solutions like MCP’s centralized metaregistry and A2A’s well-known endpoints, enterprise governance layers like Entra Agent ID, or decentralized approaches combining semantic indexing with content-addressed naming.
To move beyond research, our teams are deploying a working solution in two phases, inviting researchers and practitioners to join: Phase 1 establishes the NANDA Index at MIT as a neutral, consortium-governed switchboard for cross-organizational agent coordination; Phase 2 adds semantic discovery and federation capabilities through AGNTCY’s Agent Directory, enabling interoperability with heterogeneous registry types via DHT-based routing and OCI-aligned content addressing.
Just as ICANN enabled the internet to scale through neutral governance and open standards, realizing an Internet of AI Agents requires foundational infrastructure built collaboratively across industry, academia, and civil society, where technical standards, governance frameworks, and reference implementations reflect diverse needs rather than single-vendor control. The decisions made now will shape agent interaction patterns for decades; we invite the community to participate in building this infrastructure openly, ensuring the agentic web serves the entire ecosystem.